OpenStudy (anonymous):

Pleaaaase help met ! D: 4. Greg wants to rent a small tractor. He is trying to decide which rental company to use. Each company figures the cost on the number of hours the tractor is rented plus a delivery fee. For both companies, c represents the cost and t represents the time, in hours, of the rental. Lone Star Rentals: Canyon Farm Rentals: For how many hours of rental will the costs of both rental companies be equal? A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

were you not give the hourly rate and delivery fee for both? there is really not a way to figure this without information

OpenStudy (anonymous):

Well. You havent given the equations. But what you really have to do here is simply equate c's and solve for t.

OpenStudy (anonymous):

Pleaaaase help met ! D: 4. Greg wants to rent a small tractor. He is trying to decide which rental company to use. Each company figures the cost on the number of hours the tractor is rented plus a delivery fee. For both companies, c represents the cost and t represents the time, in hours, of the rental. Lone Star Rentals: C = 5t + 35 Canyon Farm Rentals: C = 4t + 40 For how many hours of rental will the costs of both rental companies be equal? A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

@siddhantsharan & @JuanitaM i put the equations

OpenStudy (anonymous):

set both equations to each other 5t + 35 = 4t +40 combine terms and solve for t

OpenStudy (anonymous):

5t - 4t = 40 -35 t = 5

OpenStudy (anonymous):

you good?

OpenStudy (anonymous):

Thas perfect. :D

OpenStudy (anonymous):

1. Solve: Step 1: 10x - 4 = 5x + 11 + 2x Step 2: 10x - 4 = 7x + 11 Step 3: 3x - 4 = 11 Step 4: 3x = 7 X = 7/4 Which is the first incorrect step in the solution shown above? A. Step 1 C. Step 3 B. Step 2 D. Step 4

OpenStudy (anonymous):

@JuanitaM & @siddhantsharan

OpenStudy (anonymous):

as you flow with this - what are your thoughts on the incorrect steps

OpenStudy (anonymous):

i don't know D:

OpenStudy (anonymous):

also note I can not see the original problem. Solve :

OpenStudy (anonymous):

yes i can see D as being wrong.

OpenStudy (anonymous):

D is wrong because they should have divided by 3

OpenStudy (anonymous):

Its fourth. Try anndtell mmewhy

OpenStudy (anonymous):

This is the original problem 10x - 4 = 5x + 11 + 2x

OpenStudy (anonymous):

3. Which inequality can be used to solve the following problem? Mona rented a car for $175 per week plus $0.25 per mile. How far did Mona travel if she spent at most $350 A. 175 + 0.25m < 350 B. 175 + 0.25m > 350 C. 175 + 350 < 0.25m D. 350 + 0.25m < 175

OpenStudy (anonymous):

@JuanitaM & @siddhantsharan